OpenStudy (anonymous):

A room is 24 feet long, 18 feet wide, and 9 feet high. How many square yards of wallpaper are needed to paper the four walls of the room?

OpenStudy (anonymous):

for the four walls of the room u need to calculate the lateral surface area of a cuboid....

OpenStudy (radar):

Compute perimeter 36+48=84 Walls are 9 ft high compute sq ft of paper=9x84=756 divide by 9 to get sq yds 84 sq yds,

OpenStudy (anonymous):

Lateral surface area of a cuboid = 2 (Length + Breadth) × Height now put in the values of length, breadth and height...

OpenStudy (anonymous):

@radar why calculate perimeter??????

OpenStudy (radar):

You did it too lol

OpenStudy (anonymous):

no I hv not calculated PERIMETER I am calculating area!!!!!!

OpenStudy (radar):

What do you call 2( length + breadth)

OpenStudy (anonymous):

U hv not put in the formula, and it appeared out of context, thats all......

OpenStudy (radar):

I just did it in two steps. Step 1 comput perimeter Step 2 multiply by heighth (9)

OpenStudy (anonymous):

area of paper required = 2(24 ft + 18 ft) x 9 ft = 84 ft x 9 ft = 756 ft^2

OpenStudy (radar):

and......

OpenStudy (radar):

read the question. what unit do they want the amount of wallpaper in???

OpenStudy (anonymous):

1 sqft = 1/9 sqyard or 0.111 sq yard so in sq yards the area of paper required = 756 x 0.111 = 83.92 sqyards

OpenStudy (anonymous):

or if u use 1/9, u get 84 sqyards

OpenStudy (radar):

Yes, are you could of converted all the dimensions to yds to begin with lol 6 8 3

OpenStudy (anonymous):

i guess to a student who does not know how to do the question, the explanation shud b full but simple. They hv not understood the concept or they wud not b asking :))